Kibur Plc uses a job order cost system. Overhead is applied at the rate of Br 30 per direct labor

hour. Job #251 includes a prime cost of Br 20,000 and a conversion cost of Br 25,000. For this

job, 500 direct labor hours were used. Compute the total cost of Job 251.


1
Expert's answer
2022-01-12T09:13:17-0500

Formula Total job cost = Direct materials + direct labor + applied overhead

Conversion cost = Direct labor (DL) + Applied overhead

25000=DL+15000 (30*500)

Direct Labor=10000

Prime cost = Direct materials (DM) + Direct labor

20000=DM+10000

Direct Material =10000

Total Cost=10000+10000+15000

           =Br 35,000